OpenStudy (anonymous):

a 3-mile cab ride costs $7.70. a 9-mile cab ride costs $18.50. find a linear equation that models a relationship between cost c and distance d?

OpenStudy (ghazi):

lets say these are the two points on the coordinate plane can you use coordinate geometry to find relation its easy use two point slope form equation

OpenStudy (anonymous):

Make it a pair of coordinates. (3,7.70) & (9,18.50). then find the slope: (18.5-7.7)/(9-3) (10.8)/(3) m(slope)=1.8 Now write the slope int. form of a line y=mx+b plug in the values you know to find b. (use either coordinate) 18.5=1.8(9)+b b=2.3 I hope this helps...and it's right, but that's what i got. :)

OpenStudy (ghazi):

(7.70,3) and (18.50,9) lets say cost is x and distance is Y now \[(y-9)=\frac{ 9-3 }{ 18.5-7.70 }(x-7.70)\] simplify this and get your answer

OpenStudy (anonymous):

@ghazi nice , @xXRK<3SHXx cool

OpenStudy (anonymous):

very welcome :)

OpenStudy (anonymous):

:) keep rocking

OpenStudy (anonymous):

lol, now go answer my question! haha, i'm dying in trig! lol

OpenStudy (anonymous):

where is it?

OpenStudy (anonymous):

uhm idk. I don't know how to really use this yet. lol, I bumped it to the top earlier, but i can just retype it here if you will look at it?? please :)

OpenStudy (anonymous):

go in to your question and mention me like @jasonxx i will be there